OpenStudy (anonymous):

Help me! I will become a fan and give you a medal! (:

OpenStudy (anonymous):

with what

OpenStudy (anonymous):

Hold on let me get the equation done

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

f your medal, and i dont need any more fans.

OpenStudy (anonymous):

that made heaps of sense.

OpenStudy (anonymous):

im to trilllllllllll

OpenStudy (anonymous):

really ^

OpenStudy (anonymous):

\[\frac{ \sec }{ \csc -\cot } -\frac{ \sec }{ \csc -\cot } = 2 \csc\]

ganeshie8 (ganeshie8):

\[\frac{ \sec }{ \csc -\cot } -\frac{ \sec }{ \csc -\cot } = 2 \csc \] ?

OpenStudy (anonymous):

r u serious.... oh thats better

OpenStudy (anonymous):

shut up lashaun. you a fake mf

OpenStudy (anonymous):

them fighting words...

OpenStudy (anonymous):

u dont even know me

OpenStudy (anonymous):

give him some Lashaun

OpenStudy (anonymous):

If you aren't going to help please leave

OpenStudy (anonymous):

Shaylynnwalls im really sorry

ganeshie8 (ganeshie8):

\[\frac{ \sec }{ \csc -\cot } -\frac{ \sec }{ \csc -\cot } = 2 \csc \] left hand side is becoming 0. sure the problem doesnt have any typos ?

OpenStudy (anonymous):

Its ok, it isnt your fault. Some people do not understand

OpenStudy (anonymous):

And yes i am sure.

ganeshie8 (ganeshie8):

then, its a false identity !

OpenStudy (anonymous):

Wait there is a typo. The second equation the bottom is csc theta plus cot theta

ganeshie8 (ganeshie8):

\[\frac{ \sec }{ \csc -\cot } -\frac{ \sec }{ \csc + \cot } = 2 \csc \] like this ?

OpenStudy (anonymous):

yep.

OpenStudy (anonymous):

I think you do have to change everything to cosine and sine

ganeshie8 (ganeshie8):

good idea, but guess that would complicate things

ganeshie8 (ganeshie8):

lets take the LCM of denominator and simplify a bit

OpenStudy (anonymous):

Ok, that makes more sense

ganeshie8 (ganeshie8):

\[\frac{ \sec }{ \csc -\cot } -\frac{ \sec }{ \csc + \cot } = 2 \csc \] Left Hand Side : \[\frac{ \sec (\csc + \cot) - \sec ( \csc - \cot ) }{ (\csc - \cot )(\csc + \cot ) } \]

OpenStudy (anonymous):

i will give u C===8 for free

ganeshie8 (ganeshie8):

simplify a bit

OpenStudy (anonymous):

Ok let me write this down and than I can simplify.

ganeshie8 (ganeshie8):

\[\frac{ \sec }{ \csc -\cot } -\frac{ \sec }{ \csc + \cot } = 2 \csc \] Left Hand Side : \[\frac{ \sec (\csc + \cot) - \sec ( \csc - \cot ) }{ (\csc - \cot )(\csc + \cot ) } \] \[\frac{ \sec \csc + \sec \cot - \sec \csc + \sec \cot }{ (\csc - \cot )(\csc + \cot ) } \]

OpenStudy (anonymous):

So it would be sec times everything on the top?

ganeshie8 (ganeshie8):

\[\frac{ \sec }{ \csc -\cot } -\frac{ \sec }{ \csc + \cot } = 2 \csc \] Left Hand Side : \[\frac{ \sec (\csc + \cot) - \sec ( \csc - \cot ) }{ (\csc - \cot )(\csc + \cot ) } \] \[\frac{ \sec \csc + \sec \cot) - \sec \csc + \sec \cot ) }{ (\csc - \cot )(\csc + \cot ) } \] \[\frac{ \cancel{\sec \csc} + \sec \cot) - \cancel{\sec \csc} + \sec \cot ) }{ (\csc - \cot )(\csc + \cot ) } \]

ganeshie8 (ganeshie8):

yes ! you can do that...

ganeshie8 (ganeshie8):

you do in your way on ur notes. factoring sec is a brilliant idea.. go ahead :)

ganeshie8 (ganeshie8):

\[\frac{ \sec }{ \csc -\cot } -\frac{ \sec }{ \csc + \cot } = 2 \csc \] Left Hand Side : \[\frac{ \sec (\csc + \cot) - \sec ( \csc - \cot ) }{ (\csc - \cot )(\csc + \cot ) } \] \[\frac{ \sec \csc + \sec \cot - \sec \csc + \sec \cot }{ (\csc - \cot )(\csc + \cot ) } \] \[\frac{ \cancel{\sec \csc} + \sec \cot - \cancel{\sec \csc} + \sec \cot }{ (\csc - \cot )(\csc + \cot ) } \] \[\frac{ 2 \sec \cot }{ (\csc^2 - \cot^2) } \]

OpenStudy (anonymous):

Ok so once you do that you can simplify those to cot and csc if I did that correctly

ganeshie8 (ganeshie8):

yes ! bottom equals 1 as \(\csc^2 - \cot^2 = 1\)

ganeshie8 (ganeshie8):

\[\frac{ \sec }{ \csc -\cot } -\frac{ \sec }{ \csc + \cot } = 2 \csc \] Left Hand Side : \[\frac{ \sec (\csc + \cot) - \sec ( \csc - \cot ) }{ (\csc - \cot )(\csc + \cot ) } \] \[\frac{ \sec \csc + \sec \cot - \sec \csc + \sec \cot }{ (\csc - \cot )(\csc + \cot ) } \] \[\frac{ \cancel{\sec \csc} + \sec \cot - \cancel{\sec \csc} + \sec \cot }{ (\csc - \cot )(\csc + \cot ) } \] \[\frac{ 2 \sec \cot }{ (\csc^2 - \cot^2) } \] \[2 \sec \cot\] \[2 \frac{1}{\cos} \frac{\cos}{\sin} \]

OpenStudy (anonymous):

Than you would simplify the bottom equation by taking out the cos?

ganeshie8 (ganeshie8):

exactly !

ganeshie8 (ganeshie8):

complete solution, go thru :- \(\large \frac{ \sec }{ \csc -\cot } -\frac{ \sec }{ \csc + \cot } = 2 \csc \) Left Hand Side : \(\large \frac{ \sec (\csc + \cot) - \sec ( \csc - \cot ) }{ (\csc - \cot )(\csc + \cot ) } \) \(\large \frac{ \sec \csc + \sec \cot - \sec \csc + \sec \cot }{ (\csc - \cot )(\csc + \cot ) } \) \(\large \frac{ \cancel{\sec \csc} + \sec \cot - \cancel{\sec \csc} + \sec \cot }{ (\csc - \cot )(\csc + \cot ) } \) \(\large \frac{ 2 \sec \cot }{ (\csc^2 - \cot^2) } \) \(\large 2 \sec \cot\) \(\large 2 \frac{1}{\cos} \frac{\cos}{\sin} \) \(\large 2\csc\)

OpenStudy (anonymous):

Ok thank you so much! That makes so much sense now! The final answer would be 2csc!

ganeshie8 (ganeshie8):

np :)